OpenStudy (anonymous):

Factor the quadratic x^2+12x+11

OpenStudy (mathmale):

It helps to be able to recognize what kind of trinomial you have. Yours is not a perfect square, but it does have whole number zeros/roots. When a trinomial such as this one begins with x^2, that means it will likely factor into two binomial factors: (x-a)*(x-b). Now look at that last term, the 11. What are whole number factors of 11?

OpenStudy (anonymous):

Nothing because it's prime..

OpenStudy (mathmale):

Sorry. I disagree.

OpenStudy (anonymous):

Oh, uhm... 1 and 11... Sorry aha

OpenStudy (mathmale):

Every number, including prime numbers such as 11, has 1 as a factor.

OpenStudy (mathmale):

So now you have x^2 + 12x + 11 = (x - )*(x - ). Fill in the blanks. Be sure to check your factors before telling me what you got.

OpenStudy (anonymous):

Um, x^2 + 12x + 11 = (x - 11 )*(x - 1 )

OpenStudy (mathmale):

Have you checked your factors by multiplying them together (possibly using the FOIL method)?

OpenStudy (anonymous):

Okay so what we do is we use a Square chart to check our answers such as this one, (Give me a sec to draw it)

OpenStudy (anonymous):